Home‎ > ‎Crew 133‎ > ‎

117-GreenHabReport

XO Gartrelle to Cap Comm                        

The green hab is operating normally

The fan was set to 60 degrees at 8AM
The fan is operating normally at 68 degrees at 5PM

The heater was turned down to 40 degrees F at 8AM
The heater was turned up to 75 degrees F at 5PM

The pump is operating normally.

The herb seedlings still have not sprouted. The Mars shoot is still looking healthy. 
The herb seedlings and the Mars shoot received water this morning. 

The hydroponic lettuce sprouts received water today at 9AM. All the lettuce shoots appear healthy. At 5PM the hydroponic soil appeared and felt damp. The sprouts will not be watered tomorrow. The leak in the irrigation tubes grew significantly bigger today requiring an emergency repair with duct tape. Water from the leak was elevated to a height where it 
actually reached the adjeacent hydroponic bag with ease. 

Since the repair was applied after watering effectiveness of the repair will not be known until the next watering.
 
They remaining lettuce seedlings look very healthy. They are close to out growing the root cubes. 
The pepper seedlings have not sprouted as yet. All the seedlings in root cubes received water at 8AM. All the seedlings were returned to the upper hab at 5PM. 

All the other plants received water today. 

Greenhab readings for 5PM today are as follows:

                         Current             Min                 Max
Temperature        86 Degrees F   46 Degrees F  100 Degrees F   
Humidity           13%                  10%                 19%

Meter            1.07 KwH
Duration        25.04 Hours
Overflow         0.33 Liters    Removed Today at 5PM - Checked at 8AM and at 5PM


Photos from today are attached.

This is my last report before Earth transit. It has been an honor and a privilege to work with all of the Mission Support Team. Thanks for your help 
and encouragement.

Regards,
End Transmission